OpenStudy (anonymous):

geometric sequence and, if so, find a formula that describes the sequence. can someone work out there own problem for this so I can remember how to do thanks.

OpenStudy (ikram002p):

ok lol since ratio is constant mm

OpenStudy (anonymous):

my numbers are -64 32 -16 8 -4

OpenStudy (ikram002p):

well r is the ratio -4/8=r 8/-16=r -16/32=r 32/-64=r so find r

OpenStudy (ikram002p):

a_1 is the first term which is -64

OpenStudy (ikram002p):

so it would be \(a_1 r^n\)

OpenStudy (anonymous):

Got her??